On this page, we list the main changes that affected HAWKI and this web site.

Timeline of instrument status

The following table describes the events affecting the quality or the measurement of data obtained with HAWK-I. Its content can be downloaded as a Comma Separated Values file. It is complete as of September 13th, 2020.

The Start Date column indicates when the event took place. Some events cannot be defined by a single instant, but extend over sometimes several days, weeks or months. Operation with the instrument may continue while the event occurs, in which case the Ongoing flag is set. If the event had ended, the Ongoing ? flag is lowered and End Date is given. The Type column indicates the characteristic(s) of the event: Earthquake, Maintenance, Upgrade, Element failure, Intervention, Pipeline installation at QC, Change of nominal values or thresholds, Modification of configuration file of this database, or Other. The Instrument Modes column indicates the affected mode. The column KPI indicates which quality control parameters identified as Key Performance Indicators (i.e., parameters following the most critical quantities) are affected; the column non-KPI indicates which other critical Quality Control parameters are affected. The Calibration Raw Types column indicates the raw types of the affected calibrations. The Calibration Breakpoint flag indicates if calibrations obtained before the event cannot be used for science data obtained after the event or if calibration data obtained after the event cannot be used with science data obtained before the event. The Comment column provides more details. The History column allows one to see possible past information on the entry.

The Quality Control pages for HAWK-I can be accessed through the QC overview page


  • February 29 2024, HAWK-I is back to science operation after the emergency intervention. 
  • December 12 2023,  HAWK-I is out of operation due to Filter Wheel #2 failure, an emergency intervetion has been scheduled.
  • As of September 12 2023 HAWK-I suffers from a vignetting issue in the upper right corner of Q3 with an area of 200x600 pixels. 
  • The UT4 DSM is back to science operation on the night of March 22nd in both no-AO and AO modes.
  • The UT4 DSM is back to science operation on the night of March 7th in no-AO mode. Recommissioning of the AO functionalities is planned for the week of March 14-21. Please follow the La Silla Paranal Observatory News page for future updates. 
  • Since the night of 19 Feb 2023, UT4 has been out of operation due to the failure of the deformable secondary mirror (DSM). An emergency intervention is ongoing and the facility is expected to be back on March 3rd.
  • As of 30th of August 2021, HAWK-I is back to science operation. 
  • HAWK-I is not going to be available before P108 due to the contamination issues at the entrance window and covid-19.
  • During 2017, ellipticity was detected in HAWK-I, MUSE and SINFONI observations onwards when pointing away from the wind. Problem under investigation. 
    Elongation = 100*(1 – A/B) %, where A and B are the FWHM along the major and minor axes.
    The HAWK-I data classifications are :
    A. If elongation < 10% for most stars
    B. If 10% < elongation < 20% for most stars
    C. If 20% < elongation for most stars
  • HAWK-I will not be available for operations due to the following technical activities :
    • March 29th to April 4th 2015 : installation of the 1st unit of the 4LGSF and Change of Period activities
    • April 29th to May 8th 2015 : commissioning of the 1st unit of the 4LGSF
    • July 28th to September 1st 2015 : the HAWK-I CCC maintenance, followed by the installation of GRAAL, re-commissioning of HAWK-I after the GRAAL installation and test with the 1st unit of LGSF.
    • September 21st to September 30th 2015 : UT4 maintenance. 
  • As of Dec 04, 2014, all images  affected from rounding off problems have been repaired and are available in the archive.
  • As of Oct 28, 2014, HAWKI rounding off problem is solved. 
  • Since August 05, 2014, HAWK-I is suffering of a rounding-off problem which affects the coordinates keywords (RA and DEC). We are working hard to understand and solve the problem. In the meantime these values are being corrected at the archive level.
  • HAWKI is now back and will be offered from Period 94; The filter NB1190 will also be back, and the filter NB0984 removed.
  • HAWKI has not been offered in P93. However, it will stay on UT4 Nasmyth B and carry-over programs will be regularly executed.
  • After the last recoating, UT4 M1 had some incorrect force setting, which affected HAWKI image quality. All programs observed from Sept 21, 2013 to Oct 13, 2013, might have this problem. Presently, HAWKI is undergoing an intervention to clean the entrance window, which will be completed by Oct 23. 
  • August 28, 2013: HAWK-I will be temporarily removed from the UT4 Nasmyth A focus during Period 93 and until it is necessary for AOF-related activities
  • September 2012. Important news for P90 and P91 users: the filter NB1190 has been removed to make space to the filter NB9840. P90 users who asked for NB1190 can use NB9840 instead. Later we shall inform if and when NB1190 will be back.
  • March 21 - May 3, 2012: Because of preparatory work for the installation of the Adaptive Optics Facility, Yepun is not available for operations. Observations with HAWK-I are therefore not possible.
  • Dec 01, 2010: HAWK-I underwent an intervention and won't be available until Dec 05
  • Mar 01, 2010: the fast photometry mode, suitable to study transit phenomena, is officialy offered. See the manual for continuous updates
  • Nov 23, 2009: we are implementing a fast photometry mode, suitable to study transit phenomena. The mode will be offered for Period 86
  • May 29, 2009: To be consistent with all VLT instruments, HAWKI has now a different offset scheme. Read carefully the P84 User Manual
  • May 07, 2009: A few parameters controlling the readout mode of the instrument were changed. These parameters improve the ringing seen when observing bright sources. The parameters that have been changed are stored in the headers of the images and can be used to distinguish the mode used for a given file. These parameters are DET.RSPEED changed from 6 to 8 (effectively increasing the minimum DIT from 1.3 to 1.6;DET.VOLT1.DC6 was changed from 2.4 to 2.1 Volts; and DET.VOLT1.DC7 was changed from 3.4 to 3.6 Volts. This change take effect starting the night of 2009-04-07.
  • Jan 31, 2009: The January intervention was successful, and the thermal background is now three times lower.
  • Nov 19, 2008: An intervention is planned for the beginning of January 2009 to investigate the thermal background increase.
  • Sep 1, 2008: A new HAWK-I cooling mode has raised the thermal background of the instrument to approximately 2 e/s/pix (previously it was 0.1-0.2 e/s/pix). The only filters seriously affected are NB1060,and NB1190, which will need approximately 50% more integration time for each same S/N for faint objects. The origin of this increased background is under investigation.
  • July 11, 2008: HAWK-I entrance window damage giving troubles for sky subtraction. See this report.
  • August 22, 2007: HAWK-I takes off (ESO press release)
  • August 2007: Call for proposal for Science Verification